Can someone help me on this problem?
I need to solve and check the following equation:
5x + 8 + 3x -x + 5 = 6x -3
Thanks,
Sher

Answer by ankor@dixie-net.com(22740)

The idea is to get the unknown (x) on the left and the integers on the right
We do this by adding or subtracting the same values from each side.
:
5x + 8 + 3x -x + 5 = 6x - 3
:
5x + 3x - x + 8 + 5 = 6x - 3; group like terms
:
7x + 13 = 6x - 3: did the math on like terms
:
7x - 6x = - 3 - 13; subtracted 6x and 13 from both sides
:
1x = -16
:
:
Check our solution by substituting -16 for x in the original equation
5x + 8 + 3x -x + 5 = 6x - 3
:
5(-16) + 8 + 3(-16) -(-16) + 5 = 6(-16) - 3
:
-80 + 8 - 48 + 16 + 5 = -96 - 3: remember two negatives = a plus
:
-99 = -99; checks our solution x = -16